The time it takes me to drive to my favorite store follows a uniform distribution of 20 to 30 minutes. What is the probability (decimal) that it takes me 28 minutes or less?
Answer #1

Uniform distribution between a = 20 minutes and b = 30 minutes

P(it takes 28 minutes or less) = P(X 28)

= (28 - 20)/(30 - 20)

= 8/10

= 0.8